# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

### v2

@@ Introduction
As shown in Table LABEL:tab:llm_taxonomy, the majority of LLMs employ decoder-only architectures, with T5 being a notable exception. This architectural preference stems from several inherent advantages for generative tasks, particularly the natural suitability for next-token prediction and significantly improved parameter efficiency 12. Among open-source LLMs, LLaMa 2 140 is the most popular, with 15 biomedical applications built upon it. LLaMa 3 38, Mistral 7B 141, and T5 142 are also widely used. These open-source LLMs have various sizes, from 0.06B (T5) to 671B (Deepseek R1), providing flexible options for building Biomedical RAG systems. Recently, several powerful general open-source LLMs have emerged, including, Mistral Small 3.1 40 and Gemma 3 143. Although we have not found these newer models being used for biomedical RAG systems in our survey, we believe they show promise due to their outstanding performance in general domain tasks.
For commercial LLMs, the ChatGPT series APIs maintain a dominant position, presumably attributable to their superior performance. Nevertheless, several studies have demonstrated that Claude and Gemini achieve comparable efficacy when implemented in biomedical RAG applications 85, 144, 15, thereby offering viable alternatives to ChatGPT. Furthermore, it is noteworthy that both DeepSeek and Mistral have established dual accessibility paradigms: providing open-source parameters for their base models while simultaneously offering commercial API services.
In contrast to open-source and commercial LLMs, biomedical specialized LLMs have been relatively underutilized in biomedical RAG applications. Recent studies 145, 107, 146 have demonstrated that within RAG frameworks, biomedical specialized LLMs do not consistently outperform open-source alternatives with statistical significance. Thus, the more effective methodologies for leveraging biomedical specialized LLMs within retrieval-augmented generation frameworks remain an area requiring substantial further investigation.
@@
T5 142 0.06-11B Encoder-Decoder VAIV 53, CLEAR 99, DR.KNOWS 17
ChatGLM3 147 6B Decoder-only COPD 148, ChatZOC 149
LLaMA 2 140 7-70B Decoder-only Guo et al. 27, KREIMEYER et al. 150, CaLM 26, MMRAG 151, Yu et al. 152, RAMIE 107, EyeGPT 153, KG-RAG 100, JMLR 3, BiomedRAG 4, Self-BioRAG 106, CaLM 26, Du et al. 154, SurgeryLLM 28, Kreimeyer et al. 150
LLaMA 3 38 8-70B Decoder-only MMRAG 151, Fatharanihttps et al. 155, RAMIE 107, Woo et al. 156, RAG-HPO 72, i-MedRAG 157, Hewitt et al. 144, PodGPT 34, SurgeryLLM 28
Phi-3 Mini 158 3.8B Decoder-only Fatharanihttps et al. 155
Mistral 7B 141 7B Decoder-only Boulos et al. 159, RAMIE 107, Woo et al. 156, RAGPR 80, LITURAt 62, Kreimeyer et al. 150
QWen 160 32B Decoder-only NEKO 124, Thetbanthad 127
Baichuan 161 7-13B Decoder-only ChatZOC 149
Falcon 162 7-180B Decoder-only AskFDALabel 42, CaLM 26
Zephyr 163 7B Decoder-only Moser et al. 164
AceGPT 165 7-13B Decoder-only InfectA-Chat 94
Gemma 166 1-27B Decoder-only PodGPT 34
Deepseek R1 41 671B Decoder-only Feng et al. 167
@@
ChatGPT-3.5/4/4o 35 Proprietary Decoder-only Woo et al. 156, RAGPR 80, ChatZOC 149, KG-RAG 100, Clinfo.ai 14, Yu et al. 152, ChatENT 89, CaLM 26, Rau et al. 168, Endo-chat 88, BiomedRAG 4, Lu 18, Tan 169, GNQA 77, Puts 170, DRAGON-AI 83, Choi 171, RISE 85, Du et al. 154,i-MedRAG 157 , Hewitt et al. 144, RECTIFIER 91, Kainer et al. 95, Gong et al. 49, VAIA 53, RareDxGPT 16, Selcuk et al. 94, FAVOR-GPT 87, Markey et al. 172,
Claude-3.5 36 Proprietary Decoder-only Woo et al. 156, RISE 85, Hewitt et al. 144
Gemini 37 Proprietary Multimodal Upadhyaya et al. 32, Tozuka et al. 128, MEREDITH 15
@@
MedAlpaca 173 7-13B LLaMA-based RAMIE 107
PMC-LLaMA 174 7B LLaMA-based RAMIE 107, BiomedRAG 4, Ozaki et al. 145
BioMistral 175 7B Mistral-based RAMIE 107, pRAGe 176
MEDITRON 177 7-70B LLaMA-based Ozaki et al. 145, Alkaeed et al. 146
Meerkat 178 7B Mistral-based Sohn et al. 179
@@
Domain adaptation in biomedical RAG systems constitutes a methodological framework that can be strategically implemented across various RAG pipeline stages to optimize performance for medical applications 180. This approach facilitates enhanced performance on biomedical tasks while simultaneously preserving the fundamental capabilities inherent to foundation models. Based on our systematic analysis of the literature, we propose a taxonomy of domain adaptation techniques categorized according to their implementation within specific RAG pipeline components.
@@
Retriever Stage Adaptation. Specialized retrieval components constitute the initial layer of domain adaptation, wherein techniques such as contrastive learning (implemented in Contriever 57 and MedCPT 33) enable the development of more semantically rich biomedical embeddings. These adaptations augment the system’s capacity to identify and retrieve pertinent medical documents through enhanced comprehension of domain-specific terminology and conceptual relationships. Additional retriever adaptation methodologies include continued pre-training on biomedical corpora (as demonstrated in BMRETRIEVER 55) and domain-specific architectural modifications (exemplified by LADER 97 with PubMedBERT initialization). These approaches substantially enhance retrieval precision for specialized medical literature, effectively mitigating the lexical and semantic disparities between general and biomedical language domains.
@@
Reranker Stage Adaptation. The reranking phase presents a subsequent opportunity for domain specialization, employing methodologies such as the development of dedicated biomedical rerankers (e.g., MedCPT’s CrossEnc reranker 33) and the implementation of domain-specific scoring mechanisms (e.g., BiomedRAG’s chunk scorer 4). Notably, the aforementioned Model-based rerankers and LLM-based rerankers (Sec Reranker) can be conceptualized as domain adaptation strategies operationalized within the reranking stage. These specialized rerankers, trained with domain-specific datasets and optimization objectives tailored to biomedical document ranking, significantly enhance the system’s capacity to prioritize documents based on domain-relevant criteria. Furthermore, they facilitate the identification of subtle domain-specific signals that general models typically fail to detect, thereby improving retrieval precision for specialized medical queries.
@@
LLM Stage Adaptation. At the generation stage, adaptation of the language model itself through methodologies such as fine-tuning (MEDGENIE 6, SeRTS 56), parameter-efficient approaches like Low-Rank Adaptation (LoRA) (implemented in AskFDALabel 42), and continued pre-training (RALL 27, ChatENT 89) enables more accurate synthesis and explication of biomedical information. JLMR 3 joint train LLM and retrieval model, which enhances biomedical RAG’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions. These adaptations enhance the model’s comprehension of domain knowledge, facilitate the generation of appropriate medical terminology, and help improve factual accuracy in response generation.
@@
This multi-level adaptation framework creates biomedical RAG systems that leverage the general capabilities of foundation models while incorporating the specialized knowledge representation and reasoning patterns essential for healthcare and life sciences applications. Through the strategic application of adaptation techniques at each stage of the RAG pipeline, these systems achieve enhanced precision, improved contextual understanding, and more reliable generation of biomedical information, ultimately advancing the state-of-the-art in biomedical information retrieval and synthesis.
Datasets
@@
The evaluation and development of biomedical RAG systems rely heavily on appropriate datasets. In this section, we present a comprehensive overview of datasets commonly used in biomedical RAG research, categorizing them into two main types: knowledge source datasets and medical task datasets. This categorization reflects the dual nature of RAG systems, which require both comprehensive knowledge bases and task-specific evaluation benchmarks.
@@
The effectiveness of biomedical RAG systems considerably depends on the quality, comprehensiveness, and diversity of their knowledge sources 181, 1. These knowledge repositories serve as the informational backbone from which these systems retrieve and synthesize medical content. High-quality knowledge sources can improve the generated responses, making them not only relevant but also reflect the current biomedical consensus and best practices 182, 89, 106.
Biomedical knowledge spans multiple dimensions—from basic science research to clinical guidelines, from pharmaceutical data to standardized medical terminology. To address this complexity, effective biomedical RAG systems integrate diverse knowledge sources that complement each other in scope, specialization, and format. This integration enables systems to comprehensively address the multifaceted nature of medical queries 101, 55.
Table LABEL:tab:med_knowledge shows the public knowledge source utilized in the surveyed studies. Among these sources, PubMed, UMLS 183, PMC, and Medical textbooks 184 demonstrate predominant adoption, primarily attributable to their comprehensive coverage of biomedical concepts and extensive document collections. Additionally, we notice that several studies 171, 154, 80 have implemented biomedical RAG systems based on proprietary clinical data that is unpublished due to privacy concerns, while another subset of studies 50, 185, 76, 92, 165, 186 focus on a specific medical specialty, wherein specialized medical guidelines about particular clinical specialties are leveraged for constructing a targeted knowledge source.
@@
MIMIC-IV 187 EHR Decade of hospital admissions between 2008 and 2019 with data from about 300K patients and 430K admissions Moser et al. 164
PubMed Literature Repository of over 36 million biomedical research citations and abstracts, containing about 4.5B words Clinifo.ai 14, MEDRAG 61, Self-BioRAG 106, JMLR 3, Ascle 30, RISE 85, MEREDITH 15, LITURAt 62, DR.KNOWS 17, Aftab 101, CliniqIR 52, PM-Search 24, Kreimeyer 150, Kim 53
PMC Literature 8 million full-text biomedical and life sciences articles with free access, about 13.5B words Self-BioRAG 106, Soong et al. 188, PodGPT 34
TripClick 189 Literature Biomedical literature search and retrieval dataset LADER 97
GNQA 77 Literature 3000 peer reviewed publications on on aging, dementia, Alzheimer’s and diabetes GNQA 77
UMLS 183 Knowledge Base 2 million entities for 900K biomedical concepts BiomedRAG 4, CLEAR 99, Guo 27, CliniqIR 52
ICD-10 190 Knowledge Base International statistical classification of diseases and related health problems Puts et al. 170
HPO 191 Knowledge Base Human phenotype ontology Albayrak et al. 192, RAG-HPO 72, Kainer 95
MeSH Taxonomy Hierarchical organization of biomedical and health-related topics BMRETRIEVER 55
AlzKB 125 Knowledge Graph Knowledge graph for Alzheimer’s disease research ESCARGOT 125, KRAGEN 121
DrugBank 193 Database Comprehensive database integrating detailed drug data with drug target information BMRETRIEVER 55, Kim 53
SPOKE 194 Knowledge Graph Integrating more than 40 publicly available biomedical knowledge sources of separate domains KG-RAG 100
COD 149 Knowledge Base Comprehensive ophthalmic dataset ChatZOC 149
HEALIE KG29 Knowledge Graph Drawing from various medical ontologies, resources, and insights from domain experts KAKALOU et al. 29
ADRD Caregiving 26 Corpus Collection of resources supporting Alzheimer’s disease caregivers CaLM 26
StatPearls 23 Clinical Guide Collection of 9,330 clinical decision support articles available through NCBI Bookshelf MEDRAG 61, i-MedRAG 157
Medical Textbooks 184 Educational 18 core medical textbooks commonly used in USMLE preparation MEDRAG 61, Self-BioRAG 106, JMLR 3, i-MedRAG 157
Ophthalmology textbooks 153 Educational 14 specialized ophthalmology textbooks EyeGPT 153
CheXpert 195 Image-Report 224,316 chest radiographs with associated reports CLEAR 99
@@
Medical task datasets represent essential benchmarking instruments for evaluating RAG systems within healthcare applications 2, 10. These datasets simulate real clinical information challenges, providing structured frameworks to assess how systems process, interpret, and generate medical content.
These evaluation datasets can be systematically categorized according to distinct information-handling processes in the biomedical domain: biomedical information extraction, entity recognition, question answering (QA), biomedical multiple-choice examination, dialogue, and generation tasks. Table LABEL:tab:biomedical-comprehensive presents a comprehensive overview of the evaluation datasets employed across the surveyed biomedical RAG systems. Apart from works that make use of public biomedical datasets, several research groups 81, 90, 85 in favor of human-curated questions, designing proprietary data for evaluation, and inviting healthcare professionals (e.g., physicians or nurses) to conduct assessments.
@@
ChemProt 196 2010 Information Extraction BiomedRAG 4, ChemProt
DDI 197 2013 Information Extraction BiomedRAG 4, ChemProt
NCBI 198 2014 Entity Recognition GeneGPT 199, NetMe 120, RT 74
BioASQ 200 2015 Information Retrieval BMRETRIEVER 55, SeRTS 56
MIMIC-III 20 2016 Text Summarization DR.KNOWS 17, CliniqIR 52
BC5CDR 201 2016 Entity Recognition RT 74
DC3 202 2019 Diagnostic classification CliniqIR 52
MIMIC-CXR 203 2019 Text Summarization Ascle 30
MedQuAD 204 2019 QA Ascle 30
PubMedQA 205 2019 Multiple-choice BMRETRIEVER 55, Liévin et al. 206, PodGPT 34
@@
MMLU 207 2021 Multiple-choice Self-BioRAG 106, JMLR 3, i-MedRAG 157, InfectA-Chat 94, PodGPT 34
@@
MedQA 184 2021 Multiple-choice Self-BioRAG 106, JMLR 3, Ascle 30, Liévin et al. 206, i-MedRAG 157, PodGPT 34
HealthSearchQA 208 2022 QA

### v3

@@ Introduction
As shown in Table LABEL:tab:llm_taxonomy, the majority of LLMs employ decoder-only architectures, with T5 being a notable exception. This architectural preference stems from several inherent advantages for generative tasks, particularly the natural suitability for next-token prediction and significantly improved parameter efficiency Zhao2023LLMSurvey. Among open-source LLMs, LLaMa 2 HugoTouvron2023LLaMA2 is the most popular, with 15 biomedical applications built upon it. LLaMa 3 MetaAI2024LLaMA3, Mistral 7B mistral7b_2024, and T5 2020t5 are also widely used. These open-source LLMs have various sizes, from 0.06B (T5) to 671B (Deepseek R1), providing flexible options for building Biomedical RAG systems. Recently, several powerful general open-source LLMs have emerged, including, Mistral Small 3.1 mistral_small_24b_2024 and Gemma 3 gemma-3-27b-it. Although we have not found these newer models being used for biomedical RAG systems in our survey, we believe they show promise due to their outstanding performance in general domain tasks.
For commercial LLMs, the ChatGPT series APIs maintain a dominant position, presumably attributable to their superior performance. Nevertheless, several studies have demonstrated that Claude and Gemini achieve comparable efficacy when implemented in biomedical RAG applications wang2024diabetesrag, hewitt2024llmneuropathology, Lammert2024ExpertLLM, thereby offering viable alternatives to ChatGPT. Furthermore, it is noteworthy that both DeepSeek and Mistral have established dual accessibility paradigms: providing open-source parameters for their base models while simultaneously offering commercial API services.
In contrast to open-source and commercial LLMs, biomedical specialized LLMs have been relatively underutilized in biomedical RAG applications. Recent studies ozaki2024confidence, zhan2025ramie, alkaeed2025open have demonstrated that within RAG frameworks, biomedical specialized LLMs do not consistently outperform open-source alternatives with statistical significance. Thus, the more effective methodologies for leveraging biomedical specialized LLMs within retrieval-augmented generation frameworks remain an area requiring substantial further investigation.
@@
T5 2020t5 0.06-11B Encoder-Decoder VAIV Kim2024, CLEAR lopez2025clinical, DR.KNOWS Gao2025
ChatGLM3 glm2024chatglm 6B Decoder-only COPD Copd-ChatGLM, ChatZOC luo2024retrieval
LLaMA 2 HugoTouvron2023LLaMA2 7-70B Decoder-only Guo et al. Guo2024, KREIMEYER et al. kreimeyer2024rag, CaLM parmanto2024calm, MMRAG zhan2025mmrag, Yu et al. pmlr-v225-yu23b, RAMIE zhan2025ramie, EyeGPT chen2024eyegpt, KG-RAG Soman2024, JMLR Wang2024JMLR, BiomedRAG li2025biomedrag, Self-BioRAG jeong2024improving, CaLM parmanto2024calm, Du et al. du2024cognitive, SurgeryLLM ong2024surgeryllm, Kreimeyer et al. kreimeyer2024rag
LLaMA 3 MetaAI2024LLaMA3 8-70B Decoder-only MMRAG zhan2025mmrag, Fatharanihttps et al. fatharani2025pharmacogenomics, RAMIE zhan2025ramie, Woo et al. woo2025custom, RAG-HPO garcia2024deepphenotyping, i-MedRAG xiong2025iterative, Hewitt et al. hewitt2024llmneuropathology, PodGPT jia2024podgpt, SurgeryLLM ong2024surgeryllm
Phi-3 Mini phi3mini4k_2024 3.8B Decoder-only Fatharanihttps et al. fatharani2025pharmacogenomics
Mistral 7B mistral7b_2024 7B Decoder-only Boulos et al. kamel2024chat, RAMIE zhan2025ramie, Woo et al. woo2025custom, RAGPR zheng2025integrating, LITURAt Peasley2025, Kreimeyer et al. kreimeyer2024rag
QWen qwen32b_2024 32B Decoder-only NEKO XIAO202560, Thetbanthad Thetbanthad2025
Baichuan wang2025baichuan 7-13B Decoder-only ChatZOC luo2024retrieval
Falcon almazrouei2023falconseries 7-180B Decoder-only AskFDALabel llm-regulatory-framework, CaLM parmanto2024calm
Zephyr tunstall2023zephyr 7B Decoder-only Moser et al. moser2024synthetic
AceGPT huang2024acegpt 7-13B Decoder-only InfectA-Chat selcuk2025infecta
Gemma team2024gemma 1-27B Decoder-only PodGPT jia2024podgpt
Deepseek R1 deepseekr1_2024 671B Decoder-only Feng et al. feng2025retrieval
@@
ChatGPT-3.5/4/4o OpenAI2025TextGeneration Proprietary Decoder-only Woo et al. woo2025custom, RAGPR zheng2025integrating, ChatZOC luo2024retrieval, KG-RAG Soman2024, Clinfo.ai Lozano2024ClinfoAI, Yu et al. pmlr-v225-yu23b, ChatENT long2024chatent, CaLM parmanto2024calm, Rau et al. rau2024retrieval, Endo-chat fu2024application, BiomedRAG li2025biomedrag, Lu Lu2025BoostingGPT, Tan an2025chatgpt, GNQA darnell2024biomedical, Puts puts2025icd10, DRAGON-AI toro2024dragonai, Choi Choi2024, RISE wang2024diabetesrag, Du et al. du2024cognitive,i-MedRAG xiong2025iterative , Hewitt et al. hewitt2024llmneuropathology, RECTIFIER unlu2024retrieval, Kainer et al. kainer2025effectiveness, Gong et al. Gong2024, VAIA Kim2024, RareDxGPT Zelin2024RareDisease, Selcuk et al. selcuk2025infecta, FAVOR-GPT li2024favorgpt, Markey et al. markey2025rags,
Claude-3.5 Anthropic2025GettingStarted Proprietary Decoder-only Woo et al. woo2025custom, RISE wang2024diabetesrag, Hewitt et al. hewitt2024llmneuropathology
Gemini GoogleGemini2025 Proprietary Multimodal Upadhyaya et al. Upadhyaya2024, Tozuka et al. Tozuka2024, MEREDITH Lammert2024ExpertLLM
@@
MedAlpaca han2023medalpaca 7-13B LLaMA-based RAMIE zhan2025ramie
PMC-LLaMA Wu2024PMC-LLaMA 7B LLaMA-based RAMIE zhan2025ramie, BiomedRAG li2025biomedrag, Ozaki et al. ozaki2024confidence
BioMistral labrak2024biomistral 7B Mistral-based RAMIE zhan2025ramie, pRAGe buhnila2024retrieve
MEDITRON chen2023meditron70b 7-70B LLaMA-based Ozaki et al. ozaki2024confidence, Alkaeed et al. alkaeed2025open
Meerkat kim2024small 7B Mistral-based Sohn et al. sohn2024rationale
@@
Domain adaptation in biomedical RAG systems constitutes a methodological framework that can be strategically implemented across various RAG pipeline stages to optimize performance for medical applications siriwardhana2023improving. This approach facilitates enhanced performance on biomedical tasks while simultaneously preserving the fundamental capabilities inherent to foundation models. Based on our systematic analysis of the literature, we propose a taxonomy of domain adaptation techniques categorized according to their implementation within specific RAG pipeline components.
@@
Retriever Stage Adaptation. Specialized retrieval components constitute the initial layer of domain adaptation, wherein techniques such as contrastive learning (implemented in Contriever izacard2021contriever and MedCPT jin2023medcpt) enable the development of more semantically rich biomedical embeddings. These adaptations augment the system’s capacity to identify and retrieve pertinent medical documents through enhanced comprehension of domain-specific terminology and conceptual relationships. Additional retriever adaptation methodologies include continued pre-training on biomedical corpora (as demonstrated in BMRETRIEVER xu2024bmretriever) and domain-specific architectural modifications (exemplified by LADER LADER with PubMedBERT initialization). These approaches substantially enhance retrieval precision for specialized medical literature, effectively mitigating the lexical and semantic disparities between general and biomedical language domains.
@@
Reranker Stage Adaptation. The reranking phase presents a subsequent opportunity for domain specialization, employing methodologies such as the development of dedicated biomedical rerankers (e.g., MedCPT’s CrossEnc reranker jin2023medcpt) and the implementation of domain-specific scoring mechanisms (e.g., BiomedRAG’s chunk scorer li2025biomedrag). Notably, the aforementioned Model-based rerankers and LLM-based rerankers can be conceptualized as domain adaptation strategies operationalized within the reranking stage. These specialized rerankers, trained with domain-specific datasets and optimization objectives tailored to biomedical document ranking, improve the system’s ability to prioritize documents by capturing subtle domain-specific signals that general models often fail to detect, thereby enhancing retrieval precision for specialized medical queries.
@@
LLM Stage Adaptation. At the generation stage, adaptation of the language model itself through methodologies such as fine-tuning (MEDGENIE frisoni2024generate, SeRTS Hu2024), parameter-efficient approaches like Low-Rank Adaptation (LoRA)hu2022lowrank (implemented in AskFDALabel llm-regulatory-framework), and continued pre-training (RALL Guo2024, ChatENT long2024chatent) enables more accurate synthesis and explication of biomedical information. JLMR Wang2024JMLR joint train LLM and retrieval model, which enhances biomedical RAG’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions. These adaptations enhance the model’s comprehension of domain knowledge, facilitate the generation of appropriate medical terminology, and help improve factual accuracy in response generation.
@@
This multi-level adaptation framework creates biomedical RAG systems that leverage the general capabilities of foundation models while incorporating the specialized knowledge representation and reasoning patterns required for healthcare and life sciences applications. Through the strategic application of adaptation techniques at each stage of the RAG pipeline, these systems achieve enhanced precision, improved contextual understanding, and more reliable generation of biomedical information, contributing to ongoing progress in biomedical information retrieval and synthesis.
Datasets and Evaluation
@@
The evaluation and development of biomedical RAG systems rely heavily on appropriate datasets. In this section, we present a comprehensive overview of datasets commonly used in biomedical RAG research, categorizing them into two main types: knowledge source datasets and medical task benchmarks, along with their corresponding evaluation metrics. This categorization reflects the dual nature of RAG systems, which require both comprehensive knowledge bases and task-specific evaluation benchmarks.
@@
The effectiveness of biomedical RAG systems considerably depends on the quality, comprehensiveness, and diversity of their knowledge sources gao2023retrieval, Fan2024RAG. These knowledge repositories serve as the informational backbone from which these systems retrieve and synthesize medical content. High-quality knowledge sources can improve the generated responses, making them not only relevant but also reflect the current biomedical consensus and best practices huang2024toolcalling, long2024chatent, jeong2024improving.
Biomedical knowledge spans multiple dimensions—from basic science research to clinical guidelines, from pharmaceutical data to standardized medical terminology. To address this complexity, effective biomedical RAG systems integrate diverse knowledge sources that complement each other in scope, specialization, and format. This integration enables systems to comprehensively address the multifaceted nature of medical queries Aftab2024, xu2024bmretriever.
Table LABEL:tab:med_knowledge shows the public knowledge source utilized in the surveyed studies. Among these sources, PubMed, UMLS bodenreider2004umls, PMC, and Medical textbooks jin2020disease demonstrate predominant adoption, primarily attributable to their comprehensive coverage of biomedical concepts and extensive document collections. Additionally, we notice that several studies Choi2024, du2024cognitive, zheng2025integrating have implemented biomedical RAG systems based on proprietary clinical data that is unpublished due to privacy concerns, while another subset of studies Lee2024, masanneck2025evaluating, xu2024evaluation, azimi2025evaluation, huang2024acegpt, sevgi2024medical focus on a specific medical specialty, wherein specialized medical guidelines about particular clinical specialties are leveraged for constructing a targeted knowledge source.
@@
MIMIC-IV johnson2023mimic EHR Decade of hospital admissions between 2008 and 2019 with data from about 300K patients and 430K admissions Moser et al. moser2024synthetic
PubMed Literature Repository of over 36 million biomedical research citations and abstracts, containing about 4.5B words Clinifo.ai Lozano2024ClinfoAI, MEDRAG xiong-etal-2024-benchmarking, Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, Ascle Ascle, RISE wang2024diabetesrag, MEREDITH Lammert2024ExpertLLM, LITURAt Peasley2025, DR.KNOWS Gao2025, Aftab Aftab2024, CliniqIR Abdullahi2024, PM-Search Jin2022, Kreimeyer kreimeyer2024rag, Kim Kim2024
PMC Literature 8 million full-text biomedical and life sciences articles with free access, about 13.5B words Self-BioRAG jeong2024improving, Soong et al. soong2024improving, PodGPT jia2024podgpt
TripClick rekabsaz2021tripclick Literature Biomedical literature search and retrieval dataset LADER LADER
GNQA darnell2024biomedical Literature 3000 peer reviewed publications on on aging, dementia, Alzheimer’s and diabetes GNQA darnell2024biomedical
UMLS bodenreider2004umls Knowledge Base 2 million entities for 900K biomedical concepts BiomedRAG li2025biomedrag, CLEAR lopez2025clinical, Guo Guo2024, CliniqIR Abdullahi2024
ICD-10 cms2023icd10pcs Knowledge Base International statistical classification of diseases and related health problems Puts et al. puts2025icd10
HPO jsl2024hpo Knowledge Base Human phenotype ontology Albayrak et al. albayrak2024hpo, RAG-HPO garcia2024deepphenotyping, Kainer kainer2025effectiveness
MeSH Taxonomy Hierarchical organization of biomedical and health-related topics BMRETRIEVER xu2024bmretriever
AlzKB Matsumoto2025 Knowledge Graph Knowledge graph for Alzheimer’s disease research ESCARGOT Matsumoto2025, KRAGEN Matsumoto2024
DrugBank Wishart2006 Database Comprehensive database integrating detailed drug data with drug target information BMRETRIEVER xu2024bmretriever, Kim Kim2024
SPOKE morris2023spoke Knowledge Graph Integrating more than 40 publicly available biomedical knowledge sources of separate domains KG-RAG Soman2024
COD luo2024retrieval Knowledge Base Comprehensive ophthalmic dataset ChatZOC luo2024retrieval
HEALIE KGKakalou2024 Knowledge Graph Drawing from various medical ontologies, resources, and insights from domain experts KAKALOU et al. Kakalou2024
ADRD Caregiving parmanto2024calm Corpus Collection of resources supporting Alzheimer’s disease caregivers CaLM parmanto2024calm
StatPearls Bashir2024 Clinical Guide Collection of 9,330 clinical decision support articles available through NCBI Bookshelf MEDRAG xiong-etal-2024-benchmarking, i-MedRAG xiong2025iterative
Medical Textbooks jin2020disease Educational 18 core medical textbooks commonly used in USMLE preparation MEDRAG xiong-etal-2024-benchmarking, Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, i-MedRAG xiong2025iterative
Ophthalmology textbooks chen2024eyegpt Educational 14 specialized ophthalmology textbooks EyeGPT chen2024eyegpt
CheXpert irvin2019chexpert Image-Report 224,316 chest radiographs with associated reports CLEAR lopez2025clinical
@@
Medical task datasets represent important benchmarking instruments for evaluating RAG systems within healthcare applications he2023survey, liu2024survey. These datasets simulate real clinical information challenges, providing structured frameworks to assess how systems process, interpret, and generate medical content.
These evaluation datasets can be systematically categorized according to distinct information-handling processes in the biomedical domain: biomedical information extraction, entity recognition, question answering (QA), biomedical multiple-choice examination, dialogue, and generation tasks. Table LABEL:tab:biomedical-comprehensive presents a comprehensive overview of the evaluation datasets employed across the surveyed biomedical RAG systems. Apart from works that make use of public biomedical datasets, several research groups Fukushima2025, steybe2025evaluation, wang2024diabetesrag in favor of human-curated questions, designing proprietary data for evaluation, and inviting healthcare professionals (e.g., physicians or nurses) to conduct assessments.
@@
ChemProt taboureau2010chemprot 2010 Information Extraction BiomedRAG li2025biomedrag, ChemProt
DDI segura2013semeval 2013 Information Extraction BiomedRAG li2025biomedrag, ChemProt
NCBI dogan2014ncbi 2014 Entity Recognition GeneGPT jin2024genegpt, NetMe DiMaria2024, RT li2024rt
BioASQ Tsatsaronis2015 2015 Information Retrieval BMRETRIEVER xu2024bmretriever, SeRTS Hu2024
MIMIC-III johnson2016mimic 2016 Text Summarization DR.KNOWS Gao2025, CliniqIR Abdullahi2024
BC5CDR li2016biocreative 2016 Entity Recognition RT li2024rt
DC3 eickhoff2019dc3 2019 Diagnostic classification CliniqIR Abdullahi2024
MIMIC-CXR johnson2019mimic 2019 Text Summarization Ascle Ascle
MedQuAD BenAbacha-BMC-2019 2019 QA Ascle Ascle
PubMedQA jin2019pubmedqa 2019 Multiple-choice BMRETRIEVER xu2024bmretriever, Liévin et al. Lievin2024, PodGPT jia2024podgpt
@@
MMLU hendryckstest2021 2021 Multiple-choice Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, i-MedRAG xiong2025iterative, InfectA-Chat selcuk2025infecta, PodGPT jia2024podgpt
@@
MedQA jin2020disease 2021 Multiple-choice Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, Ascle Ascle, Liévin et al. Lievin2024, i-MedRAG xiong2025iterative, PodGPT jia2024podgpt
HealthSearchQA singhal2022large 2022 QA